OpenStudy (dumbgirl211):

Is this right I'll give you a medal Combine like terms: 1. -21a + 16a A -5a B 5a C 37a D -37a** 2. 5 (x + 10) + x A 6x + 15 ** B 5x + 15 C 6x + 50 D 4x - 50

OpenStudy (danieldjpon3):

The first one is A. -5a

OpenStudy (dumbgirl211):

ok thx is the second one right

OpenStudy (danieldjpon3):

Well

OpenStudy (danieldjpon3):

What do you think is correct for question 2?

OpenStudy (dumbgirl211):

i think it's b or d

OpenStudy (danieldjpon3):

hmmm lemme think one last time

OpenStudy (dumbgirl211):

am i right

OpenStudy (danieldjpon3):

idk

OpenStudy (dumbgirl211):

ooh

OpenStudy (danieldjpon3):

I don't get the second question

OpenStudy (dumbgirl211):

hold on maybe i wrote it wrong

OpenStudy (danieldjpon3):

ye can you check please

OpenStudy (danieldjpon3):

@agent0smith Can you help us please.

OpenStudy (mathmate):

@dumbgirl211 and @danieldjpon3 It's time to review distributive law of multiplication over additon. These problems need to be evaluated to find the answers. For example, 5(4x+3)+2x =5*4x+5*3 + 2x =20x + 15 +2x =20x+2x + 15 =22x+15 If you can do the distribution, you will have no problems solving the same kind of problems.

OpenStudy (danieldjpon3):

... ._.

OpenStudy (danieldjpon3):

ok

OpenStudy (dumbgirl211):

wait i did it wrong im sorry

OpenStudy (danieldjpon3):

ty for helping us

OpenStudy (danieldjpon3):

you did?

OpenStudy (dumbgirl211):

the next question is like this

OpenStudy (danieldjpon3):

k

OpenStudy (danieldjpon3):

T_T

OpenStudy (danieldjpon3):

ok ill solve it

OpenStudy (danieldjpon3):

What you trying to find?

OpenStudy (danieldjpon3):

Oh simplify, ok

OpenStudy (danieldjpon3):

11a + 9b

OpenStudy (danieldjpon3):

is it's a variable like this 1 + b then the variable represents a 1 or a -1/0.1

OpenStudy (danieldjpon3):

Its either 11a + 11b or 11a + 9b

OpenStudy (danieldjpon3):

What do you think?

OpenStudy (dumbgirl211):

so it's the answer is b

OpenStudy (danieldjpon3):

It could be 2

OpenStudy (danieldjpon3):

what do you think

OpenStudy (dumbgirl211):

b

OpenStudy (danieldjpon3):

Ye it's B

OpenStudy (dumbgirl211):

thanks can i ask you one more

OpenStudy (danieldjpon3):

Sure!

OpenStudy (danieldjpon3):

:D

OpenStudy (danieldjpon3):

can i get my medal

OpenStudy (dumbgirl211):

its simplify

OpenStudy (danieldjpon3):

and can you open up a new question after that

OpenStudy (dumbgirl211):

ok

OpenStudy (danieldjpon3):

The same one?!

OpenStudy (dumbgirl211):

oh sorry

OpenStudy (dumbgirl211):

wait the answer to the was a or c

OpenStudy (danieldjpon3):

Which one?

OpenStudy (danieldjpon3):

this one or the previous one?

OpenStudy (dumbgirl211):

the one i just attached

OpenStudy (danieldjpon3):

ok

OpenStudy (dumbgirl211):

thank you again your helping me out so much

OpenStudy (danieldjpon3):

Maybe C

OpenStudy (danieldjpon3):

I dont know

OpenStudy (dumbgirl211):

uh ok

OpenStudy (danieldjpon3):

wait can i get the medal pls

OpenStudy (danieldjpon3):

@dumbgirl211